OpenStudy (anonymous):

What is the volume of the cylinder below? (Assume = 3.14) * object not drawn to scale 8 in 16 in Cylinder Volume = r2h A. 3,215.36 cubic inches B. 803.84 cubic inches C. 200.96 cubic inches D. 1,607.68 cubic inches

jimthompson5910 (jim_thompson5910):

The figure is missing

OpenStudy (anonymous):

oops

OpenStudy (ghazi):

volume = pi*r^2*l=3.14*(8)^2*16= 3.14*64*16=3215.36